OpenStudy (anonymous):

a car travels 420 miles in a period of 210 minutes. find the average velocity of the car in miles per hour over this period of time.

OpenStudy (anonymous):

210 minutes =210/60 hrs so in miles per hour it is 420/(210/60)

OpenStudy (anonymous):

some arithmetic here maybe \[\frac{210}{60}=\frac{7}{2}\] and \[420\div \frac{7}{2}=420\times \frac{2}{7}=120\]

OpenStudy (anonymous):

damn fast car

OpenStudy (anonymous):

hahaha thank you!

OpenStudy (anonymous):

Okay, so the way this works is as follows: We have went 420 miles. It took 210 minutes. So we need to take 420/210 to find miles/min. 420/210 = 2 Now we need to multiply this by 60 minutes to find miles/hr. So our answer is 120 miles/hr for the velocity. Hope this helps. http://www.tutorsean.net